Write the equation of the linear function g(x), for a line passing through (2,-5) and (1,3)
Answer by Alan3354(69443)

Write the equation of the linear function g(x), for a line passing through (2,-5) and (1,3)
------------
|+x +y +1|
|+2 -5 +1| = 0
|+1 +3 +1|
-----
x*(-5-3) - y*(2-1) + 1*(6+5) = 0
-8x - y = -11
8x + y = 11